OpenStudy (anonymous):

What is the value of n? 2n-7/3=15 n=?

OpenStudy (anonymous):

2n=15+7/3 n=(15+7/3)/2

OpenStudy (anonymous):

2n-7/3 is a fraction

OpenStudy (anonymous):

It doesn't matter. Get the known numbers one side, and get the unknown part the other side. Solve it as long as u know how to do addition and substraction.

OpenStudy (anonymous):

So would it be 8.6

OpenStudy (anonymous):

26/3=8.67 U are basically right.

OpenStudy (anonymous):

Ok thanks for the help